OpenStudy (anonymous):

Please show how to do this.. Find the value of sqrt(8)over 3 sqrt(4)

OpenStudy (ash2326):

Is this your question? \[\frac{\sqrt 8}{3\sqrt 4}\]

OpenStudy (anonymous):

yes:)

OpenStudy (ash2326):

Ok what are the factors of 8?

OpenStudy (anonymous):

8,16,24,32,40,48,56,64,72,80

OpenStudy (ash2326):

@ashlynnobrien those are multiples of 8, we need its factors \[8= 1\times 2 \times 4\] Isn't it?

OpenStudy (anonymous):

Ohhh, yeah. I got that confused. So your right, the factors would be 1,2,4,8

OpenStudy (ksaimouli):



OpenStudy (ash2326):

Yeah so write \(8= 2 \times 4\) and then try to solve

OpenStudy (anonymous):

That's how you solve a radical?

OpenStudy (ash2326):

Here we'll try to remove the common term from the numerator and denominator \[\frac{\sqrt 8}{3\sqrt4 }\] \[\frac{\sqrt {4 \times 2 }}{3\sqrt4 }\] \[\frac{\sqrt {4 }\sqrt{ 2 }}{3\sqrt4 }\] Can you solve now?

OpenStudy (anonymous):

you would cancel out the \[\sqrt{4}\] and be left with  right?